# AL-CoLe: Augmented Lagrangian for Constrained Learning

### v1

@@ 2.1 Statistical Constrained Learning
where $f_{\theta}(x):\mathcal{X}\to\mathcal{Y}$ is a function parametrized by the vector $\theta\subseteq\Theta\in\mathbb{R}^{p}$, $\bar{\ell}_{i}$ denotes the sample-wise loss and $\ell_{i}$ is its corresponding expected risk under distribution $\mathcal{D}_{i}$. Here $\ell_{0}$ represents the risk we wish to minimize (e.g. cross entropy) and $\ell_{i}$ the risks we wish to constrain (e.g. counterfactual fairness). For most common parametrization choices (e.g. neural networks), (P) is a non-convex constrained optimization problem.
@@ 2.2 Augmented Lagrangian and Duality
Analogous to the standard (non-augmented) dual problem, weak duality and concavity hold (see Appendix of the ArXiv version). In contrast, the augmented Lagrangian enables strong duality even in non-convex settings. We now state the assumptions required to guarantee strong duality in our problem.
@@
If Assumptions 2.12.4 hold, then strong duality holds for problem (P), i.e.,
@@
See Appendix of the ArXiv version. ∎
Theorem 2.2 establishes the existence of bounded dual variables, provided that both the first- and second-order conditions are satisfied.
@@ 2.3 Empirical Dual Learning
The dual problem (D) involves statistical functionals over unknown distributions. As customary in machine learning, we replace expected risks $\ell_{i}(f_{\theta})$ with sample means $\hat{\ell}_{i}(f_{\theta}):=\tfrac{1}{N}\sum_{n=1}^{N}\bar{\ell}_{i}(f_{\theta}(x_{n}),y_{n})$. The resulting constrained empirical risk minimization problem is
@@
Our main theoretical result is given in Theorem 2.3, which establishes bounds on both suboptimality and infeasibility of ($\hat{D}$) as an approximation of (P).
@@
Let $(\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ be a solution of the empirical augmented dual problem ($\hat{D}$). Under Assumptions 2.12.8, the solution of ($\hat{P}$) satisfies $\hat{\theta}^{\star}\in\arg\min_{\theta\in\Theta}\hat{L}(\theta,\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ and, with probability $1-(2m+1)\delta$,
$|P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})|$ $\;\leq\;(1+\Delta_{\lambda})\bar{\zeta}+\Delta_{\alpha}\bar{\zeta}^{2},$ (5)
@@
where $P^{\star}$ is the optimal value of (P), $\bar{\zeta}=\max_{i}\zeta_{i}(N_{i},\delta)$, $\Delta_{\lambda}=\max(\|\lambda^{*}\|_{1},\|\hat{\lambda}^{*}\|_{1})$, and $\Delta_{\alpha}=\max(\alpha^{*},\hat{\alpha}^{*})$.
@@
See Appendix of the ArXiv version. ∎
Inequalities (5)–(6) correspond exactly to the conditions for PACC-learnability [2, Def. 2]. Hence, Theorem 2.3 establishes that problem (P) is PACC-learnable. This is a stronger guarantee than prior results [2, 3, 4], which only achieved near-PACC-learnability [2, Def. 3], since their bound for $|P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})|$ included a constant term $\epsilon_{0}$ independent of $\bar{\zeta}$. In contrast, our augmented Lagrangian approach ensures that, with sufficient samples, the gap can be driven arbitrarily close to zero.
@@ 3 Augmented dual ascent
In the previous section we concluded that ($\hat{D}$) provides an attractive approach to obtain a PACC solution of (P), since it reduces the task to solving a concave, unconstrained problem. However, even if $(\lambda_{k},\alpha_{k})\!\to\!(\lambda^{*},\alpha^{*})$, the associated primal sequence $\{\theta_{k}^{\star}\in\arg\min_{\theta}\hat{L}(\theta,\lambda_{k},\alpha_{k})\}$ need not be asymptotically minimizing for ($\hat{P}$). Formally, an asymptotically minimizing sequence $\{\theta_{k}\}_{k=0}^{\infty}$ must satisfy:
@@
Theorem 3.1 ensures that any convergent sequence of dual iterates induces a convergent sequence of primal iterates. This is crucial in practice: any stopping point of the dual method yields a primal iterate close to optimal, without the need for averaging or randomization over iterates. To solve ($\hat{D}$), we use the increased-shifted penalty method [22, Sec 7]. The name refers to the periodic increase of $\alpha$ and corrective shifts of $\lambda$.

### v2

@@ 2.1 Statistical Constrained Learning
where $f_{\theta}(x):\mathcal{X}\to\mathcal{Y}$ is a function parametrized by the vector $\theta\subseteq\Theta\in\mathbb{R}^{p}$, $\bar{\ell}_{i}$ denotes the sample-wise loss and $\ell_{i}$ is its corresponding expected risk under distribution $\mathcal{D}_{i}$. Here $\ell_{0}$ represents the risk we wish to minimize (e.g. cross entropy) and $\ell_{i}$ the risks we wish to constrain (e.g. counterfactual fairness). For most common parametrization choices (e.g. neural networks), P is a non-convex constrained optimization problem.
@@ 2.2 Augmented Lagrangian and Duality
Analogous to the standard (non-augmented) dual problem, weak duality and concavity hold [A.1]. In contrast, the augmented Lagrangian enables strong duality even in non-convex settings. We now state the assumptions required to guarantee strong duality in our problem.
@@
If Assumptions 2.12.4 hold, then strong duality holds for problem P, i.e.,
@@
Suppose Assumptions 2.12.7 hold. Then there exists a pair $(\bar{\lambda},\bar{\alpha})$ such that
@@
See Appendix A.3. ∎
Theorem A.2 establishes the existence of bounded dual variables, provided that both the first- and second-order conditions are satisfied.
@@ 2.3 Empirical Dual Learning
The dual problem D involves statistical functionals over unknown distributions. As customary in machine learning, we replace expected risks $\ell_{i}(f_{\theta})$ with sample means $\hat{\ell}_{i}(f_{\theta}):=\tfrac{1}{N}\sum_{n=1}^{N}\bar{\ell}_{i}(f_{\theta}(x_{n}),y_{n})$. The resulting constrained empirical risk minimization problem is
@@
Our main theoretical result is given in Theorem B.1, which establishes bounds on both suboptimality and infeasibility of $\hat{D}$ as an approximation of P.
@@
Let $(\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ be a solution of the empirical augmented dual problem $\hat{D}$. Under Assumptions 2.12.8, the solution of $\hat{P}$ satisfies $\hat{\theta}^{\star}\in\arg\min_{\theta\in\Theta}\hat{L}(\theta,\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ and, with probability $1-(2m+1)\delta$,
$|P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})|$ $\;\leq\;(2+\Delta_{\lambda})\bar{\zeta}+\Delta_{\alpha}m\bar{\zeta}^{2},$ (5)
@@
where $P^{\star}$ is the optimal value of P, $\bar{\zeta}=\max_{i}\zeta_{i}(N_{i},\delta)$, $\Delta_{\lambda}=\max(\|\lambda^{*}\|_{1},\|\hat{\lambda}^{*}\|_{1})$, and $\Delta_{\alpha}=\max(\alpha^{*},\hat{\alpha}^{*})$.
@@
See Appendix B. ∎
Inequalities 1516 correspond exactly to the conditions for PACC-learnability [2, Def. 2]. Hence, Theorem B.1 establishes that problem P is PACC-learnable. This is a stronger guarantee than prior results [2, 3, 4], which only achieved near-PACC-learnability [2, Def. 3], since their bound for $|P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})|$ included a constant term $\epsilon_{0}$ independent of $\bar{\zeta}$. In contrast, our augmented Lagrangian approach ensures that, with sufficient samples, the gap can be driven arbitrarily close to zero.
@@ 3 Augmented dual ascent
In the previous section we concluded that $\hat{D}$ provides an attractive approach to obtain a PACC solution of P, since it reduces the task to solving a concave, unconstrained problem. However, even if $(\lambda_{k},\alpha_{k})\!\to\!(\lambda^{*},\alpha^{*})$, the associated primal sequence $\{\theta_{k}^{\star}\in\arg\min_{\theta}\hat{L}(\theta,\lambda_{k},\alpha_{k})\}$ need not be asymptotically minimizing for $\hat{P}$. Formally, an asymptotically minimizing sequence $\{\theta_{k}\}_{k=0}^{\infty}$ must satisfy:
@@
Theorem 3.1 ensures that any convergent sequence of dual iterates induces a convergent sequence of primal iterates. This is crucial in practice: any stopping point of the dual method yields a primal iterate close to optimal, without the need for averaging or randomization over iterates. To solve $\hat{D}$, we use the increased-shifted penalty method [22, Sec 7]. The name refers to the periodic increase of $\alpha$ and corrective shifts of $\lambda$.
@@ 5 Conclusions
Appendix A Duality of the Augmented Lagrangian
A.1 Properties of the dual problem
Proposition A.1 (Weak Duality).
For the (augmented) dual problem D, weak duality holds. Namely,
$\inf(P)\geq\sup(D)$ (12)
Proof.
Because $\Psi(x,y)\leq 0$ for all $x\leq 0$, for any feasible $\theta\in\Theta$ and any $(\lambda,\alpha)\in\mathcal{T}$ we have
$\ell_{0}(f_{\theta})\geq\ell_{0}(f_{\theta})\;+\;\alpha\sum_{i=1}^{m}\Psi(l_{i}(f_{\theta}),\tfrac{\lambda_{i}}{\alpha}).$
In particular,
$\inf(P)=\inf_{\theta\in\mathcal{F}}\ell_{0}(f_{\theta})$ $\geq\inf_{\theta\in\mathcal{F}}\Big(\ell_{0}(f_{\theta})\;+\;\alpha\sum_{i=1}^{m}\Psi(l_{i}(f_{\theta}),\tfrac{\lambda_{i}}{\alpha})\Big)$
$\geq\inf_{\theta\in\Theta}\Big(\ell_{0}(f_{\theta})\;+\;\alpha\sum_{i=1}^{m}\Psi(l_{i}(f_{\theta}),\tfrac{\lambda_{i}}{\alpha})\Big)$
$=g(\lambda,\alpha).$
Here $\mathcal{F}=\{\theta\in\Theta:\ell_{i}(f_{\theta})\leq 0,\;i=1,\dots,m\}$ denotes the feasible set, and the second inequality uses $\mathcal{F}\subseteq\Theta$. Since the bound holds for every $(\lambda,\alpha)\in\mathcal{T}$, it holds in particular at the supremum, i.e., $\sup_{(\lambda,\alpha)\in\mathcal{T}}g(\lambda,\alpha)=\sup(D)$. ∎
Proposition A.2 (Concavity of the dual).
The functions $g(\lambda,r)$ and $L(x,\lambda,r)$ are concave and upper semi-continuous for $(\lambda,r)\in T$ and non-decreasing for $r$.
Proof.
[13, Theorem 1]
A.2 Strong duality
We first recall the perturbation functions. Let $p:\mathbb{R}^{m}\to\mathbb{R}$ denote the perturbation of P:
$p(u)=$ $\min_{\theta\in\Theta}{\ell}_{0}(f_{\theta})\quad$ ($P_{u}$)
s.t. $\ell_{i}(f_{\theta})\leq u_{i},\;i\in\{1,\dots,m\}$
Likewise, $\hat{p}:\mathbb{R}^{m}\to\mathbb{R}$ is the perturbation of $\hat{P}$,
$\hat{p}(u)=$ $\min_{\theta\in\Theta}{\hat{\ell}}_{0}(f_{\theta})\quad$ ($\hat{P}_{u}$)
s.t. $\hat{\ell}_{i}(f_{\theta})\leq u_{i},\;i\in\{1,\dots,m\}$
The next two properties are the ones we show to be equivalent to strong duality.
Definition A.1 (Growth Condition).
Problem P satisfies the growth condition if there exists $\alpha>0$ such that
$L(\theta,0,\alpha)=\ell_{0}(f_{\theta})+\alpha\sum_{i=1}^{m}\max[0,\ell_{i}(f_{\theta})]^{2}$ (13)
is bounded below as a function of $\theta$.
Definition A.2 (Stability of Degree 0).
Problem P is stable of degree 0 if $p(u)$ is lower semicontinuous at $u=0$, i.e.,
$\lim_{u\downarrow 0}p(u)=p(0).$
Theorem A.1 (Strong Duality).
If Assumptions 2.12.4 hold, then strong duality holds for problem P, i.e.,
$\inf(P)=\sup(D).$
Proof.
By Assumption 2.1, $\ell_{0}(f_{\theta})$ is bounded below; let $\bar{\ell_{0}}$ be a lower bound. Since
$\alpha\sum_{i=1}^{m}\max[0,\ell_{i}(f_{\theta})]^{2}\geq 0$
for any $\alpha>0$, it follows that $L(\theta,\lambda,\alpha)$ is bounded below by $\bar{\ell_{0}}$, so the growth condition holds.
Next, by Assumption 2.3, the maps $\theta\mapsto f_{\theta}$ and $u\mapsto\bar{\ell}_{i}(u,y)$ are continuous, and $\bar{\ell}_{i}$ is uniformly bounded in $y$. Hence, by the bounded convergence theorem, $\theta\mapsto\ell_{i}(f_{\theta})$ is continuous for each $i$. Together with Assumptions 2.2 and 2.4, Berge’s Maximum Theorem applies, yielding continuity of the perturbation function $p(u)$, and in particular lower semicontinuity at $u=0$ (stability of degree 0) [25].
Finally, growth condition together with stability of degree 0 are the necessary and sufficient conditions for strong duality of the augmented dual problem D [13, Theorem 4]. ∎
A.3 Optimal Dual Solutions
Definition A.3 (Second order stability).
The problem P is stable of degree 2 if there is an open neighborhood $\mathcal{O}$ of the origin in $\mathbb{R}^{m}$, an element $\bar{\lambda}\in\mathbb{R}^{m}$ and a number $\bar{\alpha}>0$ such that
$p(u)\geq p(0)-<\bar{\lambda},u>-\frac{\bar{\alpha}}{2}||u||^{2}\quad;\quad\forall u\in\mathcal{O}$ (14)
with $p$ defined as $P_{u}$.
Theorem A.2 (Optimal Dual Solutions).
$g(\bar{\lambda},\bar{\alpha})=\max(D)=\inf(P).$
Moreover, any pair satisfies $g(\bar{\lambda},\bar{\alpha})=\max(D)$ if and only if it satisfies 14.
Proof.
By Assumption 2.1, the growth condition holds. Combined with Assumptions 2.52.7, this implies that problem P is stable of degree 2 by [13, Theorem 6]. Hence the hypotheses of [13, Theorem 5] apply, ensuring the existence of $(\bar{\lambda},\bar{\alpha})\in\mathcal{T}$ such that $g(\bar{\lambda},\bar{\alpha})=\max(D)=\inf(P)$, as well as the stated equivalence with 14. ∎
Appendix B PACC-Learnability
Theorem B.1 (PAC-Constrained Dual Learning).
Let $(\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ be a solution of the empirical augmented dual problem $\hat{D}$. Under Assumptions 2.12.8, the solution of $\hat{P}$ satisfies $\hat{\theta}^{\star}\in\arg\min_{\theta\in\Theta}\hat{L}(\theta,\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ and, with probability $1-(2m+1)\delta$,
$|P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})|$ $\;\leq\;(2+\Delta_{\lambda})\bar{\zeta}+\Delta_{\alpha}m\bar{\zeta}^{2},$ (15)
where $P^{\star}$ is the optimal value of P, $\bar{\zeta}=\max_{i}\zeta_{i}(N_{i},\delta)$, $\Delta_{\lambda}=\max(\|\lambda^{*}\|_{1},\|\hat{\lambda}^{*}\|_{1})$, and $\Delta_{\alpha}=\max(\alpha^{*},\hat{\alpha}^{*})$.
Proof.
As $(\hat{\lambda}^{\star},\hat{\alpha}^{\star})$ solves $\hat{D}$, [13, Corollary 3.2] implies that any solution of $\hat{P}$ satisfies $\hat{\theta}^{\star}\in\arg\min_{\theta\in\Theta}\hat{L}(\theta,\hat{\lambda}^{\star},\hat{\alpha}^{\star})$. It remains to establish near feasibility 16 and near optimality 15 for the statistical problem.
Near Feasibility. By Assumption 2.8, with probability $1-\delta$,
$|\,\ell_{i}(f_{\theta})-\hat{\ell}_{i}(f_{\theta})|\;\leq\;\zeta_{i}(N,\delta)$ (17)
for all $\theta$. In particular,
$\,\ell_{i}(f_{\hat{\theta^{*}}})$ $\;\leq\;\hat{\ell}_{i}(f_{\hat{\theta^{*}}})+\zeta_{i}(N,\delta)$ (18)
$\;\leq\;\zeta_{i}(N_{i},\delta)$
where the last inequality uses that $\hat{\theta}^{\star}$ is a solution (i.e., feasible) for $\hat{P}$. A union bound then gives that 18 holds simultaneously for all $i$ with probability at least $1-m\delta$.
Near Optimality. We prove the two-sided bounds
$-(2+\|\hat{\lambda}^{*}\|_{1})\bar{\zeta}-\hat{\alpha}^{*}m\bar{\zeta}^{2}$ $\;\leq\;P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})$ (19)
$P^{\star}-\ell_{0}(f_{\hat{\theta}^{\star}})$ $\;\leq\;\|{\lambda}^{*}\|_{1}\bar{\zeta}+{\alpha}^{*}m\bar{\zeta}^{2}$
where $(\lambda^{*},\alpha^{*})$ solves D, $(\hat{\lambda}^{*},\hat{\alpha}^{*})$ solves $\hat{D}$, and $\bar{\zeta}=\max_{i}\zeta_{i}(N_{i},\delta)$. Existence of these pairs follows from Assumptions 2.12.7 and Theorem A.2. Set $u_{\zeta}=\mathbf{1}_{m}\bar{\zeta}$. By near feasibility, with probability at least $1-m\delta$,
$\ell_{i}(f_{\hat{\theta^{*}}})\;\leq\;\zeta_{i}(N_{i},\delta)$ (20)
for all $i=1,\dots,m$. Hence $\hat{\theta}^{\star}$ is feasible for $P_{u}$ under the relaxation $u=u_{\zeta}$. Therefore,
$P^{*}=p(0)$ $\leq p(u_{\zeta})+<\lambda^{*},u_{\zeta}>+\alpha^{*}\|u_{\zeta}\|^{2}$ (21)
$\leq\ell_{0}(f_{\hat{\theta}^{\star}})+<\lambda^{*},u_{\zeta}>+\alpha^{*}\|u_{\zeta}\|^{2}$
$\leq\ell_{0}(f_{\hat{\theta}^{\star}})+\bar{\zeta}\|\lambda^{*}\|_{1}+\alpha^{*}m\bar{\zeta}^{2}$
with probability at least $1-m\delta$. The first inequality is by Theorem A.2; the second uses suboptimality of the feasible point $f_{\hat{\theta}^{\star}}$ for the relaxed problem.
By the analogous argument to 18, with probability $1-m\delta$,
$\,\hat{\ell}_{i}(f_{{\theta^{*}}})$ $\;\leq\;{\ell}_{i}(f_{{\theta^{*}}})+\zeta_{i}(N,\delta)$ (22)
$\;\leq\;\zeta_{i}(N_{i},\delta)$
for all $i=1,\dots,m$, where the second inequality uses feasibility of $\theta^{*}$ for P. Thus $\theta^{*}$ is feasible for $\hat{P}_{u}$ with relaxation $u=u_{\zeta}$, and so
$\hat{\ell_{0}}(f_{{\theta}^{*}})=\hat{p}(0)$ $\leq\hat{p}(u_{\zeta})+<\hat{\lambda}^{*},u_{\zeta}>+\hat{\alpha}^{*}\|u_{\zeta}\|^{2}$ (23)
$\leq\hat{\ell}_{0}(f_{{\theta}^{\star}})+<\hat{\lambda}^{*},u_{\zeta}>+\hat{\alpha}^{*}\|u_{\zeta}\|^{2}$
$\leq\hat{\ell}_{0}(f_{{\theta}^{\star}})+\bar{\zeta}\|\hat{\lambda}^{*}\|_{1}+\hat{\alpha}^{*}m\bar{\zeta}^{2}$
with probability at least $1-m\delta$. The first inequality is by Theorem A.2; the second uses suboptimality of the feasible point $f_{{\theta}^{\star}}$ for the relaxed problem. Moreover, for both $\hat{\theta}^{*}$ and $\theta^{*}$ we have
$|\hat{\ell}_{0}(f_{{\theta}})-\ell_{0}(f_{{\theta}})|\leq\zeta_{0}(N_{0},\delta)\leq\bar{\zeta}$ (24)
with probability $1-\delta$. Combining 23 and 24 and applying a union bound yields
$\ell_{0}(f_{\hat{\theta}^{*}})$ $\leq\ell_{0}(f_{\theta^{\star}})+\bar{\zeta}\|\hat{\lambda}^{*}\|_{1}+\hat{\alpha}^{*}m\bar{\zeta}^{2}+2\bar{\zeta}$ (25)
$=P^{*}+(2+\|\hat{\lambda}^{*}\|_{1})\bar{\zeta}+\hat{\alpha}^{*}m\bar{\zeta}^{2}$
with probability at least $1-(m+2)\delta$.
A final union bound shows that 21 and 25 hold simultaneously with probability $1-(2m+2)\delta$. Note that 21 already conditions on 18, so no additional union bound is required to couple near feasibility with near optimality. Taking the maximum absolute value in 19 gives 15. ∎